# Open-Sora 2.0: Training a Commercial-Level Video Generation Model in \$200k

### v2

@@ Abstract
Video generation models have achieved remarkable progress in the past year. The quality of AI video continues to improve, but at the cost of larger model size, increased data quantity, and greater demand for training compute. In this report, we present Open-Sora 2.0, a commercial-level video generation model trained for only \$200k. With this model, we demonstrate that the cost of training a top-performing video generation model is highly controllable. We detail all techniques that contribute to this efficiency breakthrough, including data curation, model architecture, training strategy, and system optimization. According to human evaluation results and VBench scores, Open-Sora 2.0 is comparable to global leading video generation models including the open-source HunyuanVideo and the closed-source Runway Gen-3 Alpha. By making Open-Sora 2.0 fully open-source, we aim to democratize access to advanced video generation technology, fostering broader innovation and creativity in content creation. All resources are publicly available at: https://github.com/hpcaitech/Open-Sora.
@@ 1 Introduction
The past year has witnessed an explosion of video generation models. Since the emergence of OpenAI’s Sora [videoworldsimulators2024] in February 2024, a series of video generation models—either open-source [zheng2024open, lin2024open, ma2025step, kong2024hunyuanvideo, hong2022cogvideo] or proprietary [polyak2024moviegen, runwayml_gen3, luma_dream_machine]—have appeared at an unprecedented pace, each striving to achieve "Sora-level" generation quality. While the quality of generated videos continues to improve, there is a clear trend toward rapid growth in model size, data quantity, and computing resources. Following the success of scaling large language models (LLMs) [hoffmann2022training, kaplan2020scaling], researchers are now applying similar scaling principles [kong2024hunyuanvideo] to video generation, converging on similar model architectures and training techniques.
In this report, however, we show that a top-performing video generation model can be trained at a highly controlled cost, offering new insights into cost-effective training and efficiency optimization.. We build Open-Sora 2.0, a commercial-level video generation model trained for only \$200k. As shown in Table. 7, the training cost is 5-10 times lower than comparable models like MovieGen [polyak2024moviegen] and Step-Video-T2V [ma2025step] based on a fair comparison. This remarkable cost efficiency stems from our joint optimization of data curation, training strategy, and AI infrastructure—all of which are detailed in the following sections.
We show the human preference evaluation of Open-Sora 2.0 and other global leading video generation models in Figure 1. The models for comparison include proprietary models like Runway Gen-3 Alpha [runwayml_gen3], Luma Ray2 [luma_dream_machine] and open-source models like HunyuanVideo [kong2024hunyuanvideo] and Step-Video-T2V [ma2025step]. We evaluate these models in three important aspects: 1) visual quality, 2) prompt adherence, and 3) motion quality. Despite the low cost of our model, it outperforms these top-performing models in at least two aspects out of the three, demonstrating its strong potential for commercial deployment.
This report is structured as follows. In Section 2, we elaborate on our data strategy, including our hierarchical data filtering system and annotation methods. Section 3 details our model architecture, covering both our novel Video DC-AE autoencoder design and the DiT architecture. Section 4 explores our cost-effective training methodology, which enables commercial-level quality at just \$200k. Section 5 presents our conditioning approaches, including image-to-video and motion control techniques. Section 6 outlines the system optimizations that maximize training efficiency, and Section 7 evaluates our model’s performance against leading video generation models. Finally, we conclude with a summary of our contributions and insights for future research.
@@ 2 Data
Our goal for data is to build a hierarchical data pyramid, catering to the requirement of the progressive training process. To this end, we develop a collection of filters that function distinctly from each other, aiming to tackle various types of data detections. By progressively strengthening the degree of filtration, we can obtain subsets of smaller sizes but higher purity and quality.
In Section 2.1, we first elaborate the data filtering system that consists of two main components: preprocessing and score filtering. Then, we introduce data annotation methods in Section 2.2. Finally, we present detailed statistics of the whole dataset in Section 2.3.
@@ 2.1 Data Filtering
The hierarchical data filtering system is illustrated in Figure 2. We begin by preprocessing raw videos into trainable video clips, then progressively apply a series of filters, ranging from loose to strict, to construct a structured data pyramid.
@@ 2.1.1 Preprocessing
The preprocessing phase converts raw videos into short clips suitable for training. During this phase, we first eliminate broken files and raw videos with outlying attributes. Specifically, we filter out videos with a duration of less than 2 seconds, a bit per pixel (bpp) below 0.02, a frame rate (fps) under 16, and an aspect ratio outside the range [1/3, 3], and those with a "Constrained Baseline" profile.
Then, we detect continuous shots within the raw video and segment it into short clips based on the detection result. For shot detection, we apply the libavfilter from FFmpeg [FFmpeg] to calculate the scene scores, which quantify visual differences among frames. The shot boundaries are identified at points where the change in scene scores exceeds an empirically determined threshold.
Finally, we segment the video based on the shot detection results, ensuring that the output clips adhere to specific format constraints: frame rate (fps) below 30, longer dimension not exceeding 1080 pixels, and H.264 codec. Additionally, black borders are removed during this process. For those shots exceeding 8 seconds, we further divide them into multiple 8-second clips, while shots shorter than 2 seconds are discarded.
@@ 2.1.2 Score Filtering
To address the various defects in raw data, we develop a bag of complimentary filters, each targeting a specific aspect of data quality. These filters work together as a comprehensive and robust purification system. Typically, each filter evaluates a sample by assigning a score based on its respective criteria, and the filtering intensity is controlled by a threshold. We introduce all the score-based filters in the following sections.
Aesthetic Score.
We assess the aesthetic quality of a sample using the CLIP+MLP aesthetic score predictor [schuhmann2021aesthetic]. For a video sample, we extract the first, middle, and last frame, compute their individual scores, and take the average as the final aesthetic score.
Motion Score.
The motion intensity of a video is measured using the VMAF motion score from FFmpeg’s libavfilter library [FFmpeg]. Videos with extremely low or excessively high motion scores are filtered out to maintain a balanced motion range.
Blur Detection.
To evaluate image clarity, we apply the Laplacian operator from OpenCV [opencv_library]. A sample is considered blurry if the variance of the Laplacian image falls below a configurable threshold. For video samples, we extract five uniformly spaced frames and adopt a majority voting approach to determine blurriness.
OCR.
We detect text bounding boxes using PaddleOCR [paddleocr2020] and compute the total area of bounding boxes with a confidence score above 0.7. Images or videos containing excessive text are discarded to avoid unwanted overlays.
Camera Jitter Detection.
We detect camera jitter using Shot Boundary Detection from the PySceneDetect library [pyscenedetect]. A video is identified as having camera jitter if the average frame-to-frame change exceeds a predefined threshold.
@@ 2.2 Data Annotation
For captioning, we utilize the open-source vision-language model LLaVA-Video [zhang2024video] to annotate 256px videos. We prompt the model to focus on six aspects for a detailed and comprehensive caption, which are 1) main subjects; 2) subjects’ actions; 3) background and environment; 4) lighting condition and atmosphere; 5) camera movement; 6) video style, such as realistic, cinematic, 3D, animation, and so on. For high-resolution 768px training data, we leverage a stronger proprietary model Qwen 2.5 Max [qwen25] to generate more accurate and semantically aligned captions. We find that Qwen 2.5 Max produces fewer hallucinations and delivers better semantic consistency than LLaVA-Video.
For both training and inference, we append the motion score after the caption to enable configurable control over motion strength in the generated videos.
2.3 Data Statistics

### v3

# arXiv:2503.09642v3[cs.GR] 02 Mar 2026

@@ Abstract
Video generation models have achieved remarkable progress in the past year. The quality of AI video continues to improve, but at the cost of larger model size, increased data quantity, and greater demand for training compute. In this report, we present Open-Sora 2.0, a commercial-level video generation model trained for only \$200k. With this model, we demonstrate that the cost of training a top-performing video generation model is highly controllable. We detail all techniques that contribute to this efficiency breakthrough, including data curation, model architecture, training strategy, and system optimization. According to human evaluation results and VBench scores, Open-Sora 2.0 is comparable to global leading video generation models including the open-source HunyuanVideo and the closed-source Runway Gen-3 Alpha. By making Open-Sora 2.0 fully open-source, we aim to democratize access to advanced video generation technology, fostering broader innovation and creativity in content creation.
Video Generation, Diffusion Transformer
@@ 1 Introduction
The past year has witnessed an explosion of video generation models. Since the emergence of OpenAI’s Sora (Brooks et al., 2024) in February 2024, a series of video generation models—either open-source (Zheng et al., 2024; Lin et al., 2024; Ma et al., 2025a; Kong et al., 2024; Hong et al., 2022) or proprietary (Polyak et al., 2024; RunwayML, 2025; Luma AI, 2025)—have appeared at an unprecedented pace, each striving to achieve ”Sora-level” generation quality. While the quality of generated videos continues to improve, there is a clear trend toward rapid growth in model size, data quantity, and computing resources. Following the success of scaling large language models (LLMs) (Hoffmann et al., 2022; Kaplan et al., 2020), researchers are now applying similar scaling principles (Kong et al., 2024) to video generation, converging on similar model architectures and training techniques.
In this paper, however, we show that a top-performing video generation model can be trained at a highly controlled cost, offering new insights into cost-effective training and efficiency optimization. We build Open-Sora 2.0, a commercial-level video generation model trained for only \$200k. As shown in Figure 6, the training cost is 5-10 times lower than comparable models like MovieGen (Polyak et al., 2024) and Step-Video-T2V (Ma et al., 2025a) based on a fair comparison. This remarkable cost efficiency stems from our joint optimization of data curation, training strategy, and AI infrastructure—all of which are detailed in the following sections.
We show the human preference evaluation of Open-Sora 2.0 and other global leading video generation models in Figure 1. The models for comparison include proprietary models like Runway Gen-3 Alpha (RunwayML, 2025), Luma Ray2 (Luma AI, 2025) and open-source models like HunyuanVideo (Kong et al., 2024) and Step-Video-T2V (Ma et al., 2025a). We evaluate these models in three important aspects: 1) visual quality, 2) prompt adherence, and 3) motion quality. Despite the low cost of our model, it outperforms these top-performing models in at least two aspects out of the three, demonstrating its strong potential for commercial deployment.
@@ 2 Data
Our goal for data is to build a hierarchical data pyramid, catering to the requirement of the progressive training process. To this end, we develop a collection of filters that function distinctly from each other, aiming to tackle various types of data detections. By progressively strengthening the degree of filtration, we can obtain subsets of smaller sizes but higher purity and quality. For completeness, we further conduct a statistical analysis of some key attributes of the collected video data, including distributions of visual and textual attributes in Appendix A.
@@ 2.1 Data Filtering
The hierarchical data filtering system is illustrated in Figure 2. We begin by preprocessing raw videos into video clips, then progressively apply a series of filters, ranging from loose to strict, to construct a structured data pyramid.
@@ 2.1.1 Preprocessing
The preprocessing phase converts raw videos into short clips suitable for training. During this phase, we first eliminate broken files and raw videos with outlying attributes. Specifically, we filter out videos with a duration of less than 2 seconds, a bit per pixel (bpp) below 0.02, a frame rate (fps) under 16, and an aspect ratio outside the range [1/3, 3], and those with a ”Constrained Baseline” profile. Then, we segment raw videos into short continuous clips based on scene scores calculated by libavfilter from FFmpeg (FFmpeg Developers, 2023). Finally, we process the obtained continuous clips, ensuring that the output clips adhere to specific format constraints: frame rate (fps) below 30, longer dimension not exceeding 1080 pixels, and H.264 codec. Additionally, black borders are removed by clipping the shots. Last, we divide shots exceeding 8 seconds into multiple 8-second clips, while shots shorter than 2 seconds are discarded.
@@ 2.1.2 Score Filtering
To address the various defects in raw data, we develop a bag of complimentary filters, including 1) aesthetic score; 2) motion score; 3) blur detection; 4) OCR; 5) camera jitter detection, each filter targets a specific aspect of data quality. These filters work together as a comprehensive and robust purification system. Typically, each filter evaluates a sample by assigning a score based on its respective criteria, and the filtering intensity is controlled by a threshold. We introduce all the score-based filters in Appendix B.
@@ 2.2 Data Annotation
For captioning, we utilize the open-source vision-language model LLaVA-Video (Zhang et al., 2024) to annotate 256px videos. We prompt the model to focus on six aspects for a detailed and comprehensive caption, which are 1) main subjects; 2) subjects’ actions; 3) background and environment; 4) lighting condition and atmosphere; 5) camera movement; 6) video style, such as realistic, cinematic, 3D, animation, and so on. For high-resolution 768px training data, we leverage a stronger proprietary model Qwen 2.5 Max (Team, 2024) to generate more accurate and semantically aligned captions. We find that Qwen 2.5 Max produces fewer hallucinations and delivers better semantic consistency than LLaVA-Video. For both training and inference, we append the motion score after the caption (detailed in Section 5.3).
3 Model Architecture
3.1 3D Autoencoder
We begin by leveraging the open-source HunyuanVideo VAE (Kong et al., 2024) as the initial autoencoder. To further reduce both training and inference costs, we develop a video autoencoder with deep compression (Video DC-AE, named after (Chen et al., 2024)) that improves efficiency while maintaining high reconstruction fidelity.
HunyuanVideo VAE achieves a compression ratio of $4\times 8\times 8$. Under that compression ratio, even with a patch size of $2\times 2$, our generation model still needs to process around 115K tokens per training video, resulting in considerable computational demands. To address the above challenge, we identify potential redundancy in the spatial dimension and propose increasing the spatial compression ratio to 32 while keeping the temporal compression ratio fixed at 4. As shown in Figure 4, during training, our Video DC-AE with a $4\times 32\times 32$ compression ratio and a patch size of $1\times 1\times 1$ enables training on 32-frame, 256px videos with a latent shape of $8\times 8\times 8$. This configuration preserves a balanced level of information across all dimensions and yields a $5.2\times$ improvement in training throughput. For generation, for instance, compared to HunyuanVideo VAE, our Video DC-AE reduces the token count from 76K to 19K for a 5-second, 24fps, 768px video. A detailed derivation of the token count reduction is provided in Appendix F.1. In addition, inference speed improves by over $10\times$, making this approach highly efficient.
Chen et al. (2024) introduce DC-AE as an effective approach to achieve higher downsampling ratios while maintaining reconstruction quality. While the original DC-AE is primarily optimized for image encoding, we extend its architecture to support video encoding by incorporating temporal compression mechanisms.
As shown in Figure 3(a), our Video DC-AE encoder consists of three residual blocks followed by three EfficientViT blocks (Cai et al., 2023), with the decoder adopting a symmetrical structure. The first five encoder blocks and the last five decoder blocks function as downsampling and upsampling layers, respectively.
To adapt DC-AE for video encoding, we replace 2D operations (e.g., convolutions, normalization) with 3D operations. Additionally, we introduce temporal compression in the last two downsampling blocks (blocks 4 and 5) of the encoder, and temporal upsampling in the first two upsampling blocks (blocks 2 and 3) of the decoder to reconstruct temporal information effectively. Moreover, we follow DC-AE to introduce special residual blocks to connect the downsample and upsample blocks. This is because Chen et al. (2024) identifies gradient propagation issues in these blocks, such that without the residuals, it is especially difficult to train high-compression autoencoders.
As shown in Figure 3(b), our downsample residual blocks follow DC-AE’s pixel-shuffling strategy, redistributing pixels from the spatial and temporal dimensions into the channel dimension, followed by channel-wise averaging to achieve the desired compression. The upsample residual blocks perform the inverse operation by first duplicating channels, then redistributing them back into the spatial and temporal dimensions to reconstruct the original structure.
Model Down. (TxHxW) Info. Down. Channel Causal LPIPS↓ PSNR↑ SSIM↑
Open-Sora 1.2 (Zheng et al., 2024) $4\times 8\times 8$ 192 4 $\checkmark$ 0.161 27.504 0.756
StepVideo VAE (Ma et al., 2025a) $8\times 16\times 16$ 96 64 $\checkmark$ 0.082 28.719 0.818
HunyuanVideo VAE (Kong et al., 2024) $4\times 8\times 8$ 48 16 $\checkmark$ 0.046 30.240 0.856
[…]
Video DC-AE $4\times 32\times 32$ 48 256 $\times$ 0.049 30.777 0.872
We train Video DC-AE from scratch and evaluate its reconstruction quality, with results summarized in Table 1. Our model is compared against Open-Sora 1.2 VAE, StepVideo VAE (Ma et al., 2025a), and HunyuanVideo VAE (Kong et al., 2024). The results demonstrate that Video DC-AE achieves competitive performance, with only minor degradation in LPIPS (Zhang et al., 2018) scores compared to the best-performing HunyuanVideo VAE, while maintaining strong PSNR and SSIM scores.
3.2 DiT Architecture
Figure 5: Open-Sora diffusion transformer architecture.
To achieve higher video quality, we employ full attention to capture long-range dependencies effectively. After encoding the input through the autoencoder, we patchify the latent representations to enhance computational efficiency and improve model learning. Sana (Xie et al., 2024) suggests that a patch size of 1 yields better training stability and finer details in video generation. With Video DC-AE’s high compression ratio, we can afford to reduce the patch size to 1 (i.e., no patch at all), whereas a patch size of 2 is still required when the HunyuanVideo autoencoder is used in our generation model.
Inspired by FLUX’s MMDiT (Labs, 2024), we employ a hybrid transformer architecture that incorporates both dual-stream and single-stream processing blocks. In the dual-stream blocks, text and video information are processed separately to facilitate more effective feature extraction within each modality. Subsequently, single-stream blocks integrate these features to facilitate effective cross-modal interactions. To further enhance the model’s ability to capture spatial and temporal information, we apply 3D RoPE (Rotary Position Embedding) (Su et al., 2024), which extends traditional positional encoding to three-dimensional space, allowing the model to better represent motion dynamics across time. For text encoding, we leverage T5-XXL (Chung et al., 2024) and CLIP-Large (Radford et al., 2021), two high-capacity pretrained models known for their strong semantic understanding. T5-XXL captures complex textual semantics, while CLIP-Large improves alignment between text and visual concepts, leading to more accurate prompt adherence in video generation.
An overview of the model architecture is illustrated in Figure 5, and the detailed architectural specifications are presented in Table 3 in Appendix C.1.
4 Model Training
A commercial-grade video generation model typically requires more than 10 billion parameters and O(100M) training samples. To mitigate the substantial computational cost associated with training, we propose a cost-effective training pipeline, as outlined in Table 2. Our approach consists of three key stages: (1) training a text-to-video (T2V) model on low-resolution video data, (2) training an image-to-video (I2V) model on low-resolution video data, and (3) fine-tuning an I2V model on high-resolution videos.
Training Stage Dataset CP #iters #GPUs #GPU day USD
256px T2V 70M 1 85k 224 2240 \$107.5k
256px T/I2V 10M 1 13k 192 384 \$18.4k
768px T/I2V 5M 4 13k 192 1536 \$73.7k
Total 4160 \$199.6k
4.1 Efficient Training Strategy
Our budget-conscious training strategy emphasizes the following four key aspects:
4.1.1 Leveraging Open-Source Image Models
Prior research (Ma et al., 2024; Hong et al., 2022; Zheng et al., 2024) has demonstrated that pretraining on image datasets can significantly accelerate video model training. To avoid the high cost of training an image model from scratch, we leverage Flux (Labs, 2024), an open-source state-of-the-art text-to-image model with 11 billion parameters, which provides sufficient capacity to generate high-quality videos. We initialize our T2V model using Flux and empirically find that, despite it being a distilled model, this initialization is effective for further training.
4.1.2 High-Quality Training Data
Inspired by the efficient image training strategies from PixArt (Chen et al., 2023), we hypothesize that high-quality video data can substantially enhance training efficiency. Consequently, we curate a high-quality subset from a large-scale dataset for low-resolution training. For high-resolution fine-tuning, we impose stricter selection criteria to ensure superior video quality.
4.1.3 Learning Motion in Low-Resolution
As shown in Table 4 and Table 5 in Appendix D, training on a 129-frame video at 768px resolution is 40 times slower than at 256px. This performance gap arises due to the quadratic computational complexity of self-attention mechanisms as the number of tokens increases. To optimize efficiency, we allocate the majority of computational resources to low-resolution training, reducing the need for expensive high-resolution computations.
4.1.4 Image-to-Video Models Facilitate Resolution Adaptation
We find that adapting a model from 256px to 768px resolution is significantly more efficient using an I2V approach compared to T2V. We hypothesize that conditioning the model on a static image allows it to focus more on motion generation, a capability that is well-learned during low-resolution training.
Based on this observation, we prioritize training an I2V model at high resolution. During inference, we first generate an image from a text prompt and subsequently synthesize a video conditioned on both the image and the text. During training, T/I2V training at low resolution followed by brief fine-tuning on high-resolution videos yields high-quality results with minimal additional training.
With our proposed approach, we successfully constrain the one-time training cost to \$200K, as detailed in Table 2. We estimate the training costs of Movie Gen (Polyak et al., 2024) and Step-Video-T2V (Ma et al., 2025a) based on publicly available information, as summarized in Figure 6. Our model achieves 5–10× lower training costs. We hope that these insights will contribute to reducing the training cost of high-quality video generation models in future research.
4.2 Training Setting
Our training setup is largely based on Open-Sora 1.2 (Zheng et al., 2024). We adopt flow matching (Lipman et al., 2022) as our primary training objective and utilize the AdamW (Loshchilov and Hutter, 2017) optimizer with $\beta$ values of (0.9, 0.999) and an $\epsilon$ value of $1\times 10^{-15}$. No weight decay is applied. The learning rate is set to $5\times 10^{-5}$ for the first 40k steps of Stages 1 and 2, followed by a decay to $3\times 10^{-5}$ for the final 45k steps. For Stage 3, it is reduced to $1\times 10^{-5}$. Gradient norm clipping is applied with a threshold of 1.
4.2.1 Training Objective
Our flow matching approach is similar to that used in Stable Diffusion 3 (Esser et al., 2024). We denote the video latent as $X_{0}$, and Gaussian noise $X_{1}\sim\mathcal{N}(0,1)$. The model $f_{\theta}$ takes as input an interpolated latent $X_{t}=(1-t)X_{0}+tX_{1}$ and is trained to predict the velocity component $X_{0}-X_{1}$. The corresponding loss function is formulated as:
where $y$ represents the conditioning input (text and/or image). The timestep $t$ is first sampled from a logit-normal distribution and then scaled according to the shape of $X_{0}$. Given that higher-resolution and longer-duration videos are more susceptible to noise, we apply the following transformation: $t\leftarrow\frac{\alpha t}{1+(\alpha-1)t}$ during training and inference, where $\alpha$ is proportional to the product $T\times H\times W$.
4.2.2 Multi-Bucket Training
Following the methodology in Open-Sora 1.2 (Zheng et al., 2024), we adopt multi-bucket training to efficiently handle videos of varying frame counts, resolutions, and aspect ratios within the same batch. This strategy optimizes GPU utilization by dynamically assigning batch sizes based on video characteristics. We conduct a search on H200 GPUs to determine batch sizes used for different training configurations. The specific batch sizes and hyperparameter searching methods are detailed in Appendix D.
4.3 High Compression Autoencoder Adaptation
4.3.1 Challenges in Training Video Autoencoders.
With the growing interest in video generation models, we observe that despite variations in architecture, downsampling ratios, number of channels, and training losses, video autoencoders trained with similar information downsampling ratios exhibit comparable performance (evaluated at 256px resolution) (Chen et al., 2024). We define the information downsampling ratio, $D_{\text{info}}$ as follows:
However, training a deep video compression network is challenging. Prior work (Chen et al., 2024) identifies gradient propagation issues in downsample and upsample blocks. To address this, they introduce residual connections in these blocks using pixel shuffle and unshuffle operations. We verify that training the HunyuanVideo VAE architecture with residual connections improves performance. Building on this, we incorporate temporal downsampling into DC-AE to construct Video DC-AE.
[…]
4.3.2 Challenges in Training the Generation Model.
While training throughput for the diffusion model improves, we find that high-compression autoencoders—especially those with larger latent channel dimensions—slow down convergence in training the diffusion model. Prior studies (Yao and Wang, 2025; Xie et al., 2024) show that when training image diffusion models on AEs with better reconstruction but higher latent dimensionality, the generation quality often degrades. Additionally, Theorem A.7 in (Chen et al., 2025) suggests that, under strong assumptions, increasing the channel size by $k$ requires $k^{5}$ more data to achieve comparable performance. Meanwhile, Yao and Wang (2025) demonstrates that introducing a distillation loss between a pretrained image foundation model and the VAE latents can significantly accelerate diffusion training.
4.3.3 Strategies for Using High-Compression Autoencoders.
  1. 1.
    Latent Space Distillation: After training Video DC-AE, we apply a distillation loss to align the third-layer latents with DINOv2 (Oquab et al., 2023), leveraging their similar latent shapes.
  2. 2.
  3. 3.
    Prioritizing I2V Training: We observe that I2V models adapt more efficiently than T2V models when switching to high-compression AEs. Thus, we focus on training an I2V model for this setting.
  4. 4.
    Tiling in video encoding: High compression AEs trained at low resolutions experience performance degradation when reconstructing high-resolution videos (Chen et al., 2024). Although we could have fine-tuned the Video DC-AE at high resolutions, we re-use the tiling code by Kong et al. (2024) to save training resources.
We first train the model on short videos (up to 33 frames) for 17K iterations using 20M samples across 160 GPUs. Then, we extend training to long videos (up to 128 frames) for 8K iterations on 2M samples. The final model achieves a loss of 0.5, compared to 0.1 for the untrained model. However, due to computational constraints, training does not fully converge. As shown in Figure 13 in Appendix E, while the fast video generation model underperforms the original, it still captures spatial-temporal relationships. We release this model to the research community for further exploration.
4.3.4 Discussion and Future Work.
We believe high-compression autoencoders are critical for end-to-end video generation, especially for high-resolution videos that contain substantial redundancy. To further reduce training costs, it is essential to: (1) Optimize autoencoder training to yield better latent spaces for diffusion-based learning. (2) Design high-throughput autoencoders, as autoencoder encoding time remains a major computational bottleneck in generative models using high-compression autoencoders.
4.4 AE Training
$\mathcal{L}=\mathcal{L}_{1}+0.5\mathcal{L}_{\text{LPIPS}},$
5 Conditioning
5.1 Image-to-Video Training
Figure 7: Image and Video Condition Framework. Condition information is injected into the model via an additional channel dimension. A mask mechanism is introduced to distinguish different input types. This design enables support for a wide range of I2V and video-to-video (V2V) tasks.
As illustrated in Figure 7, our approach first encodes image or video conditions using an autoencoder, then concatenates the encoded features with the original video latent representation. An extra channel is introduced to indicate the task type, increasing the number of channels from $k$ to $2k+1$ .
To improve generalization, we introduce image condition dropout similar to text condition dropout. During training, dropping the image condition reduces the problem to a T2V setting, where zero tensors are concatenated to the video latent. For T/I2V training, we set the dropout ratio to 12.5%, ensuring robustness across various I2V and T2V tasks.
5.2 Image-to-Video Inference
Figure 8: Heatmap of Image Guidance Scale Across Denoising Steps and Latent Frames. Darker regions indicate higher image guidance values, emphasizing stronger influence on later frames and earlier denoising steps.
We use classifier-free guidance for inference (Ho and Salimans, 2022). As our model is conditioned on both image and text, a straightforward approach is to use a single guidance scale:
$v_{t}$ $=v_{\theta}(x_{t},t,\varnothing,\varnothing)$
A further issue arises when using high image guidance, as it sometimes introduces flickering in generated videos. To mitigate this, we introduce a guidance oscillation technique (Ho et al., 2022) for image guidance. Taking 50-step sampling as an example, after the first 10 steps, we alternate the image guidance scale: during odd-numbered steps, we apply the original $g_{\textit{img}}$, while for even-numbered steps, we reduce it to 1. This helps balance stability and motion consistency.
5.3 Motion Score
Controlling the level of motion strength is a crucial feature in video generation. We explicitly model the motion dynamics as a separate controllable parameter. We leverage the motion score obtained from data pre-processing, which quantifies the level of dynamics in a video, by appending it to the caption as an additional conditioning signal. During inference, adjusting the motion score allows for effective and independent control over the dynamic level of the generated video, as shown in Figure 9.
5.4 Inference-Time Scaling
We additionally explore an inference-time scaling strategy inspired by Ma et al. (2025b) to enhance video generation quality and diversity without modifying the underlying model. Since this technique is not used in our main evaluation or comparative benchmarking, we defer the implementation details, and qualitative results to Appendix G.
6 System Optimization
We train our models using ColossalAI (Li et al., 2023), an efficient parallel training system. Our hardware setup includes H200 GPUs, whose 141GB memory enables more effective data parallelism (DP) and allows for more aggressive selective activation checkpointing, significantly optimizing resource utilization. Additionally, we leverage PyTorch compile (Ansel et al., 2024) and Triton kernels (OpenAI, 2019) to accelerate training efficiency.
To efficiently handle high-resolution video training, we employ multiple parallelization techniques. For video autoencoders, we adapt tensor parallelism (TP) (Shoeybi et al., 2020) to convolution layers by partitioning weights either the input or output channel dimensions, reducing memory consumption while preventing out-of-bound indexing for high-resolution training. For MMDiT training, we combine Zero Redundancy Optimizer (ZeroDP) (Rajbhandari et al., 2020) with Context Parallelism (CP) (Liu et al., 2023). Here, video and text sequences are partitioned across GPUs along the sequence dimension, enabling each GPU to compute attention independently. This approach mitigates the memory bottleneck and improves attention computation efficiency, particularly for high-resolution videos where attention complexity grows quadratically.
Empirical evaluations on H200 GPUs (141GB memory) indicate that employing CP alone yields an optimal trade-off between memory efficiency and computational performance. During Stage 1 and Stage 2 training, we exclusively utilize DP in combination with ZeRO-2, achieving a maximum FLOPs utilization (MFU) of 38.19%. In Stage 3, we integrate ZeRO-2 with CP=4, resulting in a MFU of 35.75%.
Beyond parallelization, our training system incorporates several established optimization techniques, including selective activation checkpointing, automatic failure recovery, using pinned-memory buffers for the dataloader, and the saving and loading of checkpoints. We defer detailed descriptions of these auxiliary optimizations to Appendix H.
7 Performance
Our model supports both T2V and I2V generation at 256×256px and 768×768px resolutions (hereafter referred to as 256px and 768px), generating videos up to 128 frames long. At 24 FPS, this corresponds to a 5-second duration. Since our model is optimized for I2V generation, the default text-to-image-to-video (T2I2V) pipeline first generates an image using the FLUX model, which is then used as the starting frame for video generation. The generated results are presented in Figure 16 in Appendix I.
We further evaluate our model’s performance using VBench (Huang et al., 2024) in Figure 10, demonstrating significant improvements from Open-Sora 1.2 to 2.0. The performance gap between Open-Sora and OpenAI’s Sora has been reduced from 4.52% to 0.69%, highlighting substantial advancements in video generation quality. Additionally, our model achieves a higher VBench score compared to CogVideoX1.5-5B and HunyuanVideo, further establishing its superiority among current open-source T2V models.
Figure 10: VBench Score Comparison. Our model outperforms open-source T2V models by leveraging a text-to-image-to-video generation approach. Additionally, our latest version significantly narrows the performance gap between Open-Sora and OpenAI’s Sora, demonstrating substantial improvements in video generation quality and coherence.
8 Conclusion
This paper introduces Open-Sora 2.0, a commercial-level video generation model that was trained for only \$200k, which is 5-10 times more cost-efficient than comparable models like MovieGen and Step-Video-T2V. This achievement highlights that high-quality video generation models can be developed with highly controlled costs through careful optimization of data curation, model architecture, training strategy, and system optimization. Despite its significantly lower training cost, Open-Sora 2.0 performs comparably to leading video generation models including HunyuanVideo and Runway Gen-3 Alpha. The model supports both T2V and I2V generation at resolutions up to 768×768 pixels for videos up to 5 seconds in length.
Looking ahead, several challenges remain in video generation. First, deep compression video VAE technology is still underexplored. While we aggressively increase the compression ratio to reduce latent tokens, this approach introduces reconstruction quality loss and adaptation difficulties. Additionally, diffusion models often produce artifacts such as object distortion and unnatural physics, with users having limited control over these details. The field needs further research on artifact prevention and enhanced control over generated content. We hope that by open-sourcing Open-Sora 2.0, we can provide the community with tools to collectively tackle these challenges, fostering innovation and advancements in the field of video generation.
9 Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Data Statistics